# From Exploration to Mastery: Enabling LLMs to Master Tools via Self-Driven Interactions

### v1

@@ Abstract
Tool learning enables Large Language Models (LLMs) to interact with external environments by invoking tools, serving as an effective strategy to mitigate the limitations inherent in their pre-training data. In this process, tool documentation plays a crucial role by providing usage instructions for LLMs, thereby facilitating effective tool utilization. This paper concentrates on the critical challenge of bridging the comprehension gap between LLMs and external tools due to the inadequacies and inaccuracies inherent in existing human-centric tool documentation. We propose a novel framework, DRAFT, aimed at Dynamically Refining tool documentation through the Analysis of Feedback and Trails emanating from LLMs’ interactions with external tools. This methodology pivots on an innovative trial-and-error approach, consisting of three distinct learning phases: experience gathering, learning from experience, and documentation rewriting, to iteratively enhance the tool documentation. This process is further optimized by implementing a diversity-promoting exploration strategy to ensure explorative diversity and a tool-adaptive termination mechanism to prevent overfitting while enhancing efficiency. Extensive experiments on multiple datasets demonstrate that DRAFT’s iterative, feedback-based refinement significantly ameliorates documentation quality, fostering a deeper comprehension and more effective utilization of tools by LLMs. Notably, our analysis reveals that the tool documentation refined via our approach demonstrates robust cross-model generalization capabilities.
@@ 1 Introduction
Figure 1: (a): An illustration showcasing issues in current tool documentation, including incomplete, redundant, and inaccurate information. (b): An illustration of how LLMs incrementally enhance their understanding of tool usage through iterative exploration, transitioning from untested functionalities (gray) to discovered capabilities (white) via a trial-and-error process. (c): The comparison between the raw and improved tool documentation regarding its usefulness on three tool learning datesets, highlighting that DRAFT is able to produce tool documentations that are more favored by LLMs.
@@
Humans, on the contrary, acquire proficiency in tool usage through repeated interactions and hands-on experiences, capable of maintaining an updated comprehension of these tools despite their evolving functionalities. In light of this, this paper proposes a novel framework, DRAFT, conceptualized to automate the adjustment and optimization of tool documentation based on the outcomes and feedback derived from the LLM’s interaction with the tool, aiming explicitly at bridging the comprehension gap between LLMs and external tools.
More concretely, DRAFT implements a trial-and-error methodology to incrementally improve the tool documentation. As shown in Figure 2, DRAFT orchestrates three dynamically interlinked phases: experience gathering, learning from experience, and documentation rewriting, which collectively facilitate the iterative process of documentation enhancement. It first undertakes the simulation of potential tool application scenarios, crafting explorative instances and capturing tool execution outcomes through a designed explorer. Subsequently, the analyzer dissects the prevailing documentation, amalgamating insights from the explorer’s findings and feedback to moot documentation modification propositions. Finally, the rewriter amalgamates these insights, refining the tool documentation while simultaneously guiding further explorative pursuits by the explorer. To optimize this process, we design a diversity-promoting exploration strategy to ensure diversity in exploration, thus providing a wider range of samples for subsequent rewriting. Recognizing that different tools vary in complexity for LLMs to comprehend, we introduce a tool-adaptive termination mechanism to improve efficiency during modifications by halting the iterative process once the documentation aligns with the comprehension of LLMs, thereby saving time and resources while preventing overfitting. Through this collaborative, iterative framework, the documentation progressively morphs, aligning more coherently with LLMs operational requisites, and thus empowering LLMs to leverage external tools more effectively in their problem-solving endeavors.
@@
Figure 2: Schematic illustration of our proposed self-driven iterative improvement framework, DRAFT. The left part illustrates the three distinct learning phases in our framework. The middle part shows the two specialized mechanisms designed to optimize the process. The right part depicts an example of using DRAFT to iteratively modify the tool documentation.
@@ 2.1 Overview of DRAFT
To address the challenges of inadequate, ambiguous, and outdated tool documentation that hinder LLMs from effectively utilizing external tools, we propose DRAFT, a framework that iteratively refines tool documentation to bridge the comprehension gap between LLMs and tools. As illustrated in Figure 2, DRAFT operates through three interconnected phases: experience gathering, learning from experience, and documentation rewriting. In the experience gathering phase, an explorer simulates diverse tool usage scenarios, collecting data on how the LLM interacts with the tool based on the current documentation, thus uncovering misunderstandings and limitations. The learning from experience phase involves an analyzer examining this data to identify discrepancies between intended and actual tool usage, pinpointing ambiguities or inaccuracies in the documentation, and proposing targeted improvements. In the documentation rewriting phase, a rewriter integrates these insights to update the documentation, enhancing clarity and alignment with the tool’s functionalities. Through this trial-and-error framework, DRAFT is capable of simulating the process by which humans acquire proficiency in tool usage through repeated interactions and hands-on experiences, thereby automating the creation of tool documentation specifically designed for LLMs. Furthermore, by employing the diversity-promoting exploration strategy and tool-adaptive termination mechanism, DRAFT efficiently converges on optimized documentation, enabling LLMs to utilize external tools more effectively despite the initial documentation shortcomings.
@@ 2.2 Experience Gathering
In the experience gathering phase, we design an Explorer $\mathcal{A}_{E}$ to simulate plausible scenarios in which the tool may be utilized. This approach parallels the manner in which individuals investigate the potential applications of a new tool when they are unable to comprehend the accompanying manual.
Specifically, at the $i$-th iteration, the Explorer generates an exploration instance $e_{i}$ based on the current tool documentation $t_{i-1}$, next-step exploration direction $d_{i-1}$ from the Rewriter, and the previous history $\mathcal{H}_{i}=\{(e_{j},r_{j})|j<i\}$, which consists of prior exploration instances $e_{<i}$ and their corresponding return results of the tool $r_{<i}$. This process is formalized as follows:
@@
Given that tool utilization often involves complex parameter ranges, combinations, and potential error sources, it is crucial to ensure diversity in the exploration phase to cover a wide spectrum of possible scenarios. To address this, besides maintaining a record of all previously explored queries—which we provide to the Explorer to instruct it to generate instances that differ from those already generated—we also implement a diversity-promoting exploration strategy:
@@
Self-Reflection. If the similarity constraint is not satisfied (i.e., the new query is too similar to previous ones), the Explorer engages in self-reflection. It discards the current instance and analyzes the reasons for the overlap, adjusting its approach to generate a new query that explores different aspects of the tool. This iterative self-reflection process continues until the Explorer produces an instance that meets the diversity criterion.
@@ 2.3 Learning from Experience
Analogous to how humans learn—acquiring familiarity with new tools through practical experiences and then consulting manuals to deepen understanding—the insights gained during the experience gathering phase provide a foundation for informed and targeted enhancements to the documentation. Thus, building upon the experience gathered in the first phase, the second phase focuses on analyzing this data to refine the tool documentation. In this phase, we introduce an Analyzer designed to identify and address issues within the current tool documentation, thereby guiding the Rewriter in making effective revisions.
Formally, at the $i$-th iteration, the Analyzer takes the following inputs: current tool documentation $t_{i-1}$, exploration instance $e_{i}$, tool feedback $r_{i}$ provided by the Explorer, and the history of documentation revisions $\mathcal{T}_{i}=\{t_{j}\mid j<i\}$. Then the Analyzer analyzes these inputs to identify issues and generate revision suggestions $s_{i}$:
@@
To ensure that the Analyzer can provide high-quality and relevant revision suggestions, we establish several evaluation criteria, including consistency with tool outputs, comprehensiveness, and conciseness without irrelevant information (as detailed in Appendix B). These criteria enable Analyzer to identify and assess issues present within the existing tool documentation. By considering the historical evolution of the documentation through the revision history $\mathcal{T}_{i}$, the Analyzer gains valuable insights into past updates, helping it avoid redundant or repetitive suggestions and focus on areas that still require improvement. The prompt template for Analyzer is illustrated in Figure 4.
@@ 2.4 Documentation Rewriting
Specifically, we design a Rewriter to synthesize information from the exploration instances $e_{i}$ and the corresponding tool return results $r_{i}$ provided by the Explorer, as well as the revision suggestions $s_{i}$ from the Analyzer. It is important to note that the Rewriter also takes into account the rewrite history $\mathcal{T}_{i}$, which includes all previous versions of the tool documentation up to iteration $i$. By integrating these inputs, the Rewriter produces an updated version of the tool documentation $t_{i}$ and provides suggestions for the next round of exploration directions $d_{i}$. This process is formalized as:
@@
Furthermore, recognizing that different tools vary in their complexity and the ease with which LLMs can comprehend them, we implement a tool-adaptive termination mechanism to adaptively determine when to cease modifications for each tool. Analogous to recipes requiring different levels of expertise, some tools may reach optimal documentation faster than others. We consider the iterative process to have converged when there is minimal change between two consecutive versions of the documentation, indicating that the Rewriter has sufficiently aligned the documentation with the LLM’s understanding.
Specifically, we measure the degree of change $\Delta$ between iterations by calculating both the word-match metric (e.g., BLEU score (Papineni et al., 2002)) and the semantic-match metric (e.g., cosine similarity of embeddings):
@@
where $\mathbf{e}^{t}_{i}$ and $\mathbf{e}^{t}_{i-1}$ are the embedding vectors of $t_{i}$ and $t_{i-1}$ obtained using OpenAI’s text-embedding-ada-002. The function $\mathrm{sim}(\cdot,\cdot)$ calculates the cosine similarity between the semantic embedding vectors of two documentation versions, and $\mathrm{BLEU}(\cdot,\cdot)$ measures the n-gram overlap between them. If $\Delta$ exceeds a predefined termination threshold $\tau$, we stop the iterative modifications.
@@ 3.1 Experimental Setup
Implementation Details. For the main experiments, we use the GPT-4o as the backbone model for DRAFT, which means that we employ this model to refine the tool documentation by incorporating its own tool usage feedback. We set the similarity threshold $\phi$ to $0.9$ for diversity-promoting exploration strategy, and termination threshold $\tau$ to $0.75$ for tool-adaptive termination mechanism. We select three of the latest LLMs to ensure our evaluation reflects the current state of the field including the closed-source models GPT-4o  [^2] https://platform.openai.com/playground/chat?models=gpt-4o-2024-08-06 and GPT-4o-mini  [^3] https://platform.openai.com/playground/chat?models=gpt-4o-mini, as well as the open-source model Llama-3-70B  [^4] https://huggingface.co/meta-llama/Meta-Llama-3-70B. Our code is available at https://github.com/quchangle1/DRAFT.
@@ 3.2 Experimental Results
Table 1: Performance comparison of different methods on three datasets. CP%, Winare short for Correct Path Rate and Win Rate, respectively. Winis calculated by comparing each method with ReAct. The term “-” means that EayTool has not been implemented on the Spotify dataset. The best result for each LLM is in bold.
@@ 3.3 Further Analysis
How does the number of iteration rounds affect the performance of tool learning? A key feature of DRAFT is its iterative refinement of tool documentation with a tool-adaptive termination mechanism. We examine the effectiveness of these designs by analyzing how the number of iteration rounds influences the performance of downstream tool learning. The results presented in Figure 6 demonstrate a general trend where performance enhances with an increasing number of iterations, followed by a subsequent decline. This suggests that iterative modifications are crucial for enhancing tool documentation quality and the ability of LLMs to utilize tools effectively. Such modifications facilitate the exploration of a wider array of examples and the incorporation of additional feedback derived from the tool usage experiences of LLMs, ultimately refining the tool descriptions to achieve superior performance. However, a decline in performance is observed after a certain number of iterations. This decline may be attributed to the introduction of redundant information as the number of iterations increases, potentially leading to overfitting. Therefore, we implement a tool-adaptive termination mechanism to prevent performance degradation and ensure optimal results.
@@
Does using other models as backbones also ensure cross-model generalization? Our preliminary experiments have demonstrated that employing GPT-4o as the backbone to refine tool documentation by integrating its own usage feedback results in revised documentation that exhibits cross-model generalization, enhancing the performance of other models. This observation prompts an inquiry into whether using other models as a backbone would similarly yield cross-model generalization. To investigate this, we conduct experiments on RestBench-TMDB using Llama-3-70B as a backbone, which led to the generation of a new set of tool documentation informed by its usage feedback. As illustrated in Figure 7, our findings indicate that it also demonstrates cross-model generalization, enhancing the tool usage capabilities across all LLMs. This may result from decoder-only models sharing similar transformer structures and common pre-training corpora, allowing them to achieve consensus on knowledge mastery and needs. Furthermore, we also find that employing GPT-4o as a backbone yields superior performance compared to LLama-3-70B. This suggests that our method can benefit from the continuous improvements in foundational models.
@@
Are the two mechanisms we proposed truly effective for DRAFT? We conduct ablation studies on RestBench-TMDB using GPT-4o to assess the impact of the two mechanisms incorporated within DRAFT. The results presented in Table 2, highlight the significance of each mechanism: w/o diversity refers to a variant that the explorer generates exploration instances without applying similarity constraint, meaning that all generated instances are considered. The absence of the diversity-promoting exploration strategy leads to a notable performance drop, underscoring the need for its implementation to enhance exploration diversity. w/o adaptive refers to a variant that does not terminate prematurely based on the iteration conditions; rather, it guarantees that each iteration of the tool documentation completes the designated maximum number of iteration rounds before ceasing operations. The significant decline in performance observed in this variant further emphasizes the benefits of introducing the tool-adaptive termination mechanism, which serves to prevent overfitting while improving efficiency. The results with other LLMs and datasets demonstrate similar trends and are provided in Appendix A.
@@
Does the tool documentation modified by DRAFT also improve human comprehension? In the previous experiments, we have demonstrated that tool documentation modified by DRAFT can help LLMs better understand and utilize external tools. Next, we aim to verify whether this improvement extends to human understanding of the tools as well. Specifically, we conduct a human evaluation, inviting three well-educated doctor students to evaluate the raw tool documentation and the documentation modified by DRAFT based on three criteria: completeness: whether the documentation is comprehensive, conciseness: whether it is clear and concise, accuracy: whether it reflects the tool’s functionality accurately. We randomly sampled 50 cases from RestBench and ToolBench for this evaluation.
@@
The results, as shown in Table LABEL:tab:quality_evaluation, indicate that the tool documentation modified by DRAFT demonstrates significant improvements, particularly in terms of completeness and accuracy (some cases are presented in Appendix C). This can be attributed to the fact that DRAFT incorporates feedback from LLMs based on their tool usage experience, which provides execution results during the revision process, ensuring better alignment between the documentation and the actual tool behavior. Moreover, our design of a tool-adaptive termination mechanism effectively prevents unnecessary redundancy, leading to more concise documentation. Additionally, the results suggest that ToolBench tools have significantly lower documentation quality than those in the RestBench dataset, as the readability of ToolBench tools exhibited a more substantial improvement after modification by DRAFT. This further highlights the necessity of revising the existing tool documentation.

### v2

@@ Abstract
Tool learning enables Large Language Models (LLMs) to interact with external environments by invoking tools, serving as an effective strategy to mitigate the limitations inherent in their pre-training data. In this process, tool documentation plays a crucial role by providing usage instructions for LLMs, thereby facilitating effective tool utilization. This paper concentrates on the critical challenge of bridging the comprehension gap between LLMs and external tools due to the inadequacies and inaccuracies inherent in existing human-centric tool documentation. We propose a novel framework, DRAFT, aimed at Dynamically Refining tool documentation through the Analysis of Feedback and Trials emanating from LLMs’ interactions with external tools. This methodology pivots on an innovative trial-and-error approach, consisting of three distinct learning phases: experience gathering, learning from experience, and documentation rewriting, to iteratively enhance the tool documentation. This process is further optimized by implementing a diversity-promoting exploration strategy to ensure explorative diversity and a tool-adaptive termination mechanism to prevent overfitting while enhancing efficiency. Extensive experiments on multiple datasets demonstrate that DRAFT’s iterative, feedback-based refinement significantly ameliorates documentation quality, fostering a deeper comprehension and more effective utilization of tools by LLMs. Notably, our analysis reveals that the tool documentation refined via our approach demonstrates robust cross-model generalization capabilities.
@@ 1 Introduction
Figure 1: (a): An illustration showcasing issues in current tool documentation, including incomplete, redundant, and inaccurate information. (b): An illustration of how LLMs incrementally enhance their understanding of tool usage through iterative exploration, transitioning from untested functionalities (gray) to discovered capabilities (white) via a trial-and-error process. (c): The comparison between the raw and improved tool documentation regarding its usefulness on three tool learning datasets, highlighting that DRAFT is able to produce tool documentation that is more favored by LLMs.
@@
Humans, on the contrary, acquire tool proficiency through repeated interactions and hands-on experiences, capable of maintaining an updated comprehension of these tools despite their evolving functionalities. In light of this, this paper proposes DRAFT, conceptualized to automate the adjustment and optimization of tool documentation based on the feedback derived from the LLM’s interaction with the tool, aiming to bridge the comprehension gap between LLMs and external tools.
More concretely, DRAFT implements a trial-and-error methodology to incrementally improve the tool documentation. As shown in Figure 2, DRAFT orchestrates three dynamically interlinked phases, which collectively facilitate the iterative process of documentation enhancement. It first undertakes the simulation of potential tool application scenarios, crafting explorative instances and capturing tool execution outcomes through a designed explorer. Subsequently, the analyzer dissects the prevailing documentation, amalgamating insights from the explorer’s findings and feedback to moot documentation modification propositions. Finally, the rewriter amalgamates these insights, refining the tool documentation while simultaneously guiding further explorative pursuits by the explorer. To optimize this process, we design a diversity-promoting exploration strategy to ensure diversity in exploration, thus providing a wider range of samples for subsequent rewriting. Recognizing that different tools vary in complexity for LLMs, we introduce a tool-adaptive termination mechanism to improve efficiency during modifications by halting the iterative process once the documentation aligns with the comprehension of LLMs, thereby saving time and resources while preventing overfitting. Through this collaborative, iterative framework, the documentation progressively morphs, aligning more coherently with LLMs operational requisites, and thus empowering LLMs to leverage external tools more effectively in their problem-solving endeavors.
@@
Figure 2: Schematic illustration of our proposed self-driven iterative improvement framework, DRAFT. The left part illustrates the three distinct learning phases in our framework, where a trial refers to a single iteration of the process. The middle part shows the two specialized mechanisms designed to optimize the process. The right part depicts an example of using DRAFT to iteratively modify the tool documentation (a complete revision trajectory is presented in Appendix E).
@@ 2.1 Overview of DRAFT
To address the challenges of inadequate, ambiguous, and outdated tool documentation that hinder LLMs from effectively utilizing external tools, we propose DRAFT, a framework that iteratively refines tool documentation to bridge the comprehension gap between LLMs and tools. As illustrated in Figure 2, DRAFT operates through three interconnected phases: experience gathering, learning from experience, and documentation rewriting. In the experience gathering phase, an explorer simulates diverse tool usage scenarios, collecting data on how the LLM interacts with the tool based on the current documentation, thus uncovering misunderstandings and limitations. The learning from experience phase involves an analyzer examining this data to identify discrepancies between intended and actual tool usage, pinpoint ambiguities or inaccuracies in the documentation, and propose targeted improvements. In the documentation rewriting phase, a rewriter integrates these insights to update the documentation, enhancing clarity and alignment with the tool’s functionalities. Through this trial-and-error framework, DRAFT is capable of simulating the process by which humans acquire proficiency in tool usage through repeated interactions and hands-on experiences, thereby automating the creation of tool documentation specifically designed for LLMs. Furthermore, by employing the diversity-promoting exploration strategy and tool-adaptive termination mechanism, DRAFT efficiently converges on optimized documentation, enabling LLMs to utilize external tools more effectively despite the initial documentation shortcomings.
@@ 2.2 Experience Gathering
In the experience gathering phase, we design an Explorer $\mathcal{M}_{E}$ to simulate plausible scenarios in which the tool may be utilized. This approach parallels the manner in which individuals investigate the potential applications of a new tool when they are unable to comprehend the accompanying manual.
Specifically, at the $i$-th iteration, the Explorer $\mathcal{M}_{E}$ generates an exploration instance $e_{i}$ based on the current tool documentation $t_{i-1}$, next-step exploration direction $d_{i-1}$ from the Rewriter $\mathcal{M}_{R}$, and the previous history $\mathcal{H}_{i}=\{(e_{j},r_{j})|j<i\}$, which consists of prior exploration instances $e_{<i}$ and their corresponding return results of the tool $r_{<i}$. This process is formalized as follows:
@@
Given that tool utilization often involves complex parameter ranges, combinations, and potential error sources, it is crucial to ensure diversity in the exploration phase to cover a wide spectrum of possible scenarios (Hong et al., 2018; Friedrich et al., 2009). To address this, besides maintaining a record of all previously explored queries—which we provide to the Explorer to instruct it to generate instances that differ from those already generated—we also implement a diversity-promoting exploration strategy:
@@
Self-Reflection. If the similarity constraint is not satisfied (i.e., the new query is too similar to previous ones), the Explorer engages in self-reflection (Shinn et al., 2024). It discards the current instance and analyzes the reasons for the overlap, adjusting its approach to generate a new query that explores different aspects of the tool. This iterative self-reflection process continues until the Explorer produces an instance that meets the diversity criterion.
@@ 2.3 Learning from Experience
Analogous to how humans learn—acquiring familiarity with new tools through practical experiences and then consulting manuals to deepen understanding—the insights gained during the experience gathering phase provide a foundation for informed and targeted enhancements to the documentation. Thus, building upon the experience gathered in the first phase, the second phase focuses on analyzing this data to refine the tool documentation. In this phase, we introduce an Analyzer $\mathcal{M}_{A}$ designed to identify and address issues within the current tool documentation, thereby guiding the Rewriter in making effective revisions.
Formally, at the $i$-th iteration, the Analyzer $\mathcal{M}_{A}$ takes the following inputs: current tool documentation $t_{i-1}$, exploration instance $e_{i}$, tool feedback $r_{i}$ provided by the Explorer $\mathcal{M}_{E}$, and the history of documentation revisions $\mathcal{T}_{i}=\{t_{j}\mid j<i\}$. Then the Analyzer $\mathcal{M}_{A}$ analyzes these inputs to identify issues and generate revision suggestions $s_{i}$:
@@
To ensure that the Analyzer can provide high-quality and relevant revision suggestions, we establish several evaluation criteria, including consistency with tool outputs, comprehensiveness, and conciseness without irrelevant information (as detailed in Appendix C). These criteria enable Analyzer to identify and assess issues present within the existing tool documentation. By considering the historical evolution of the documentation through the revision history $\mathcal{T}_{i}$, the Analyzer gains valuable insights into past updates, helping it avoid redundant or repetitive suggestions and focus on areas that still require improvement. The prompt template for Analyzer is illustrated in Figure 4.
@@ 2.4 Documentation Rewriting
Specifically, we design a Rewriter $\mathcal{M}_{R}$ to synthesize information from the exploration instances $e_{i}$ and the corresponding tool return results $r_{i}$ provided by the Explorer $\mathcal{M}_{E}$, as well as the revision suggestions $s_{i}$ from the Analyzer $\mathcal{M}_{A}$. It is important to note that the Rewriter $\mathcal{M}_{R}$ also takes into account the rewrite history $\mathcal{T}_{i}$, which includes all previous versions of the tool documentation up to iteration $i$. By integrating these inputs, the Rewriter $\mathcal{M}_{R}$ produces an updated version of the tool documentation $t_{i}$ and provides suggestions for the next round of exploration directions $d_{i}$. This process is formalized as:
@@
Furthermore, recognizing that different tools vary in their complexity and the ease with which LLMs can comprehend them (Qin et al., 2023b; Osiurak & Heinke, 2018), we implement a tool-adaptive termination mechanism to adaptively determine when to cease modifications for each tool. Analogous to recipes requiring different levels of expertise, some tools may reach optimal documentation faster than others. We consider the iterative process to have converged when there is minimal change between two consecutive versions of the documentation, indicating that the Rewriter has sufficiently aligned the documentation with the LLM’s understanding.
Specifically, inspired by Wieting et al. (2019), we measure the degree of change $\Delta$ between iterations by calculating both the word-match metric (e.g., BLEU score (Papineni et al., 2002)) and the semantic-match metric (e.g., cosine similarity of embeddings):
@@
where $\mathbf{e}^{t}_{i}$ and $\mathbf{e}^{t}_{i-1}$ are the embedding vectors of $t_{i}$ and $t_{i-1}$ obtained using OpenAI’s text-embedding-ada-002222https://openai.com/index/new-and-improved-embedding-model/. The function $\mathrm{sim}(\cdot,\cdot)$ calculates the cosine similarity between the semantic embedding vectors of two documentation versions, and $\mathrm{BLEU}(\cdot,\cdot)$ measures the n-gram overlap between them. If $\Delta$ exceeds a predefined termination threshold $\tau$, we stop the iterative modifications.
@@ 3.1 Experimental Setup
Implementation Details. For the main experiments, we use the GPT-4o as the backbone model for DRAFT, which means that we employ this model to refine the tool documentation by incorporating its own tool usage feedback. We set the similarity threshold $\phi$ to $0.9$, termination threshold $\tau$ to $0.75$, and maximum iteration count to $5$. We select three of the latest LLMs to ensure our evaluation reflects the current state of the field including the closed-source models GPT-4o  [^3] https://platform.openai.com/playground/chat?models=gpt-4o-2024-08-06 and GPT-4o-mini  [^4] https://platform.openai.com/playground/chat?models=gpt-4o-mini, as well as the open-source model Llama-3-70B  [^5] https://huggingface.co/meta-llama/Meta-Llama-3-70B. Our code is available at https://github.com/quchangle1/DRAFT.
@@ 3.2 Experimental Results
Table 1: Performance comparison of different methods on three datasets. Winis calculated by comparing each method with ReAct. The term “-” means that EasyTool has not been implemented on the Spotify dataset. The best result for each LLM is in bold.
@@ 3.3 Further Analysis
How does the number of iteration rounds affect the performance of tool learning? A key feature of DRAFT is its iterative refinement of tool documentation with a tool-adaptive termination mechanism. We examine the effectiveness of these designs by analyzing how the number of iteration rounds influences the performance of downstream tool learning. The results presented in Figure 6 demonstrate a general trend where performance is enhanced with an increasing number of iterations, followed by a subsequent decline. This suggests that iterative modifications are crucial for enhancing tool documentation quality and the ability of LLMs to utilize tools effectively. Such modifications facilitate the exploration of a wider array of examples and the incorporation of additional feedback derived from the tool usage experiences of LLMs, ultimately refining the tool descriptions to achieve superior performance. However, a decline in performance is observed after a certain number of iterations. This decline may be attributed to the introduction of redundant information as the number of iterations increases, potentially leading to overfitting. Therefore, we implement a tool-adaptive termination mechanism to prevent performance degradation and ensure optimal results.
@@
Does using other models as backbones also ensure cross-model generalization? Our preliminary experiments have demonstrated that employing GPT-4o as the backbone to refine tool documentation by integrating its own usage feedback results in revised documentation that exhibits cross-model generalization, enhancing the performance of other models. This observation prompts an inquiry into whether using other models as a backbone would similarly yield cross-model generalization. To investigate this, we conduct experiments on RestBench-TMDB using Llama-3-70B as a backbone, which led to the generation of a new set of tool documentation informed by its usage feedback. As illustrated in Figure 7, our findings indicate that it also demonstrates cross-model generalization, enhancing the tool usage capabilities across all LLMs. This may result from decoder-only models sharing similar transformer structures and common pre-training corpora, allowing them to achieve consensus on knowledge mastery and needs. Furthermore, we also find that employing GPT-4o as a backbone yields superior performance compared to LLama-3-70B. This suggests that our method can benefit from continuous improvements in foundational models.
@@
Are the two mechanisms we proposed truly effective for DRAFT? We conduct ablation studies on RestBench-TMDB using GPT-4o to assess the impact of the two mechanisms incorporated within DRAFT. The results presented in Table 2, highlight the significance of each mechanism: w/o diversity refers to a variant that the explorer generates exploration instances without applying similarity constraint, meaning that all generated instances are considered. The absence of the diversity-promoting exploration strategy leads to a notable performance drop, underscoring the need for its implementation to enhance exploration diversity. w/o adaptive refers to a variant that does not terminate prematurely based on the iteration conditions; rather, it guarantees that each iteration of the tool documentation completes the designated maximum number of iteration rounds before ceasing operations. The significant decline in performance observed in this variant further emphasizes the benefits of introducing the tool-adaptive termination mechanism, which serves to prevent overfitting while improving efficiency. The results with other LLMs and datasets demonstrate similar trends and are provided in Appendix B.
@@
Does the tool documentation modified by DRAFT also improve human comprehension? In the previous experiments, we have demonstrated that tool documentation modified by DRAFT can help LLMs better understand and utilize external tools. Next, we aim to verify whether this improvement extends to human understanding of the tools as well. Specifically, we conduct a human evaluation, inviting three well-educated doctor students to evaluate the raw tool documentation and the documentation modified by DRAFT based on three criteria: completeness: which documentation is more comprehensive, conciseness: which documentation is clearer and more concise, accuracy: which documentation reflects the tool’s functionality more accurately (details are presented in Appendix A). We randomly sampled 50 cases from RestBench and ToolBench for this evaluation.
@@
The results, as shown in Table LABEL:tab:quality_evaluation, indicate that the tool documentation modified by DRAFT demonstrates significant improvements, particularly in terms of completeness and accuracy (some cases are presented in Appendix D). This can be attributed to the fact that DRAFT incorporates feedback from LLMs based on their tool usage experience, which provides execution results during the revision process, ensuring better alignment between the documentation and the actual tool behavior. Moreover, our design of a tool-adaptive termination mechanism effectively prevents unnecessary redundancy, leading to more concise documentation. Additionally, the results suggest that ToolBench tools have significantly lower documentation quality than those in the RestBench dataset, as the readability of ToolBench tools exhibited a more substantial improvement after modification by DRAFT. This further highlights the necessity of revising the existing tool documentation.